The kinetic energy of a car is 8 ´ 106 J as it travels along a horizontal road. How much power is required to stop the car in 10 s?


1
Expert's answer
2022-06-07T08:11:48-0400
"P=\frac{W}{t}=\frac{\Delta E}{t}=\frac{8*10^6}{10}=8*10^5\:\rm W"